# ToolGen: Unified Tool Retrieval and Calling via Generation

### v2

@@ Abstract
As large language models (LLMs) advance, their inability to autonomously execute tasks by directly interacting with external tools remains a critical limitation. Traditional methods rely on inputting tool descriptions as context, which is constrained by context length and requires separate, often inefficient, retrieval mechanisms. We introduce ToolGen, a paradigm shift that integrates tool knowledge directly into the LLM’s parameters by representing each tool as a unique token. This enables the LLM to generate tool calls and arguments as part of its next token prediction capabilities, seamlessly blending tool invocation with language generation. Our framework allows the LLM to access and utilize a vast amount of tools with no additional retrieval step, significantly enhancing both performance and scalability. Experimental results with over 47,000 tools show that ToolGen not only achieves superior results in both tool retrieval and autonomous task completion but also sets the stage for a new era of AI agents that can adapt to tools across diverse domains. By fundamentally transforming tool retrieval into a generative process, ToolGen paves the way for more versatile, efficient, and autonomous AI systems. ToolGen enables end-to-end tool learning and opens opportunities for integration with other advanced techniques such as chain-of-thought and reinforcement learning, thereby expanding the practical capabilities of LLMs [^1] Code and data are available at: https://github.com/Reason-Wang/ToolGen.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities as interactive systems, adept at processing external inputs, executing actions, and autonomously completing tasks (Gravitas, 2023; Qin et al., 2023; Yao et al., 2023; Shinn et al., 2023; Wu et al., 2024a; Liu et al., 2024). Among the various methods enabling LLMs to interact with the world, tool calling via APIs has emerged as one of the most common and effective approaches. However, as the number of tools grows into the tens of thousands, existing methods for tool retrieval and execution struggle to scale efficiently.
@@ 2.1 Tool Retrieval
Recently, generative retrieval has emerged as a promising new paradigm, wherein models directly generate relevant document identifiers rather than relying on traditional retrieval mechanisms (Wang et al., 2022; Sun et al., 2023; Kishore et al., 2023; Mehta et al., 2023; Chen et al., 2023b). Motivated by this, ToolGen represents each tool as a unique token, allowing tool retrieval and calling to be framed as a generation task. Beyond simplifying retrieval, this design integrates smoothly with other LLM and LLM-based agent features like chain-of-thought reasoning (Wei et al., 2023) and ReAct (Yao et al., 2023). By consolidating retrieval and task execution into a single LLM agent, it reduces latency and computational overhead, leading to more efficient and effective task completion.
@@ 2.2 LLM-Agents with Tool Calling
LLMs have shown strong potential in mastering tools for various tasks. However, most existing works focus on a limited set of actions (Chen et al., 2023a; Zeng et al., 2023; Yin et al., 2024; Wang et al., 2024). For instance, Toolformer (Schick et al., 2023) fine-tunes GPT-J to handle just five tools, such as calculators. While effective for narrow tasks, this approach struggles in real-world scenarios with vast action spaces. ToolBench (Qin et al., 2023) expands the scope by introducing over 16,000 tools, highlighting the challenge of tool selection in complex environments.
To perform tool selection, current methods often use a retriever-generator pipeline, where relevant tools are retrieved and then utilized by the LLM (Patil et al., 2023; Qin et al., 2023). However, pipelined approaches face two major issues: error propagation from the retrieval step and the inability of LLMs to fully understand and use tools via simple prompting.
@@
Our approach differs from ToolkenGPT in several ways. First, we focus on real-world tools that require flexible parameters for complex tasks (e.g., YouTube channel search), while ToolkenGPT is limited to simpler tools with fewer inputs (e.g., math functions with two numbers). Additionally, ToolkenGPT relies on few-shot prompting, whereas ToolGen incorporates tool knowledge directly into the LLM through full-parameter fine-tuning, enabling the model to retrieve and execute tasks autonomously. Finally, our experiments involve a much larger tool set—47,000 tools compared to ToolkenGPT’s 13–300.
@@ 3.6 Inference
During inference, the LLM may generate action tokens outside the predefined tool token set. To prevent this, we designed a constrained beam search generation that restricts the output tokens to the tool token set. We applied this constrained beam search for both tool retrieval, where the model selects tools based on queries, and the end-to-end agent system, significantly reducing hallucination during the action generation step. A detailed analysis can be found in Section 5.4. The implementation details can be found in Appendix C.
@@ 4.1 Experimental Setup
Our experiments are based on ToolBench, a real-world tool benchmark containing more 16k tool collections, each containing several APIs, resulting in a total of 47k unique APIs. Each API is documented with a dictionary, containing the name, description, and parameters for calling the API. A real example is shown in Appendix A. We take each API as an action and map it to a token. Our retrieval and end-to-end agent-tuning data are converted from the original data in ToolBench. Details can be found in Appendix G. Although each tool may consist of multiple APIs, for simplicity, we refer to each API as a tool in this paper.
@@ 4.3 Indexing Method Comparison
Figure 3: The distribution of the number of subtokens per tool varies across different indexing methods. Atomic indexing ensures that each tool is a single token, while numeric indexing encodes tools into N tokens for tools numbered in $(10^{N-1},10^{N}]$. In contrast, both semantic indexing and hierarchical indexing produce a variable number of subtokens, with semantic indexing having more outliers with significantly longer sequences.
The implementation details are described in Appendix B.
First, we conducted an analysis of the number of subtokens required to represent each tool for the different methods, as shown in Figure 3. The figure highlights the superiority of atomic indexing, where each tool is represented by a single token, whereas other methods require multiple tokens. This efficiency allows ToolGen to reduce the number of generation tokens and inference time in both the retrieval and agent scenarios.
Next, we examined the effectiveness of different indexing methods. As shown in Table 2, semantic indexing demonstrates the best retrieval performance across various metrics and scenarios, while atomic indexing closely follows in many cases. We attribute this to the fact that semantic indexing aligns better with the pretraining data of LLMs. However, this advantage diminishes as the training data and type increase. For example, in Section 5.3, we show that atomic indexing achieves better end-to-end results. Taking all these factors into account, we choose atomic indexing for ToolGen tool virtualization.
@@ 4.4 Ablation
We perform an ablation study to assess the impact of different training stages of ToolGen, as shown in Table 3. The results indicate that retrieval training is the crucial factor for tool retrieval performance, as it directly aligns with the retrieval task where inputs are queries and outputs are tool tokens. Removing tool memorization leads to a minor performance drop although it plays a role in improving generalization, which we will discuss further in Appendix F. Similarly, constrained beam search, while not a major contributor to retrieval task, helps prevent hallucinations, making it useful for end-to-end agent tasks, see Section 5.4.
@@ 5.1 Experimental Setup
We make several modifications to the trajectory data from ToolBench to fit it into ToolGen framework. For example, as ToolGen does not require explicit selection of related tools as input, we remove this information in the system prompt. Further details are provided in Appendix G. Following this, we fine-tune the retrieval model using the reformatted data, resulting in an end-to-end ToolGen agent.
@@
  • GPT-3.5: We use gpt-3.5-turbo-0613 as one of our baselines. The implementation is the same as used in StableToolBench (Guo et al., 2024), where the tool calling capability of GPT-3.5 is used to form a tool agent.
  • ToolLlama-2: Qin et al. (2023) introduced ToolLlama-2 by fine-tuning Llama-2 (Touvron et al., 2023) model on ToolBench data.
  • ToolLlama-3: To ensure a fair comparison, we fine-tuned Llama-3, the same base model used in ToolGen, on the ToolBench dataset, creating the ToolLlama-3 baseline. In the rest of this paper, we refer to ToolLlama-3 as ToolLlama to distinguish it from ToolLlama-2.
@@
  • w/ Ground Truth Tools (G.T.) Following Qin et al. (2023), we define ground truth tools for a query as those selected by ChatGPT. For ToolLlama, we directly input the ground truth tools in the prompt, consistent with its training data format. For ToolGen, which is not trained on data with pre-selected tools, we add a prefix during the planning phase: I am using the following tools: [tool tokens], where [tool tokens] are virtual tokens corresponding to the ground-truth tools.
  • w/ Retriever In the end-to-end experiments, we use a retrieval-based setting. For baselines, we use the tools retrieved by ToolRetriever as the relevant tools. In contrast, ToolGen generates tool tokens directly, so no retriever is used.
@@
Additionally, we introduce a retry mechanism for all models to prevent early termination, the details are introduced in Section D. Specifically, if a model generates a response containing give up or I’m sorry, we prompt the model to regenerate the response with a higher temperature.
@@ 5.2 Results
Table 4 presents the end-to-end evaluation performance of various models in two settings: using Ground Truth Tools (G.T.) and a Retriever (R.). In the G.T. setting, ToolGen achieves the best average SoPR score of 54.19, outperforming GPT-3.5 and ToolLlama, with SoWR also highest for ToolGen at 49.70. In the Retriever setting, ToolGen maintains its lead with an average SoPR of 53.28 and SoWR of 51.51. ToolLlama shows competitive performance, surpassing ToolGen on some individual instances. An ablation study of end-to-end ToolGen is provided in Appendix G
@@ 6 Conclusions
Appendix A Real Tool Example
@@
Appendix B Tool Virtualization Implementation
@@
indexing is the method we use in ToolGen. Compared to other methods, it takes a single token as a tool and does not hallucinate to nonexistent tools. We use <<tool name&&api name>> to combine the tool name and api name to form a single token. For example, for the example in Appendix A, the resulting token is <<Youtube Hub&&Get Video Details>>.
@@
indexing maps each tool to the name used in ToolBench, which is also a combination between tool name and API name. However, the name can be tokenized into multiple tokens so that the model can perceive its semantic meanings. For the example in Appendix A, the resulted mapping is get_video_details_for_youtube_hub.
@@
indexing maps each tool to a unique number. We first get a list of all tools, with a length about 47,000. For all tools, we use a five digit number separated by space to represent the tool. If the example in Appendix A is the 128th element in the list, we use 0 0 0 1 2 8 to represent the tool. Since Llama-3 tokenizer encodes each number
@@
also maps each tool into a number. Different from Numeric indexing, we inject structure information into the tool representation by iterative clustering. During each iteration, we cluster tools into ten clusters, where each cluster is assigned a number from 0 to 9. For each cluster, we repeat this clustering process until there is only one tool in the cluster. These steps form a clustering tree. We take the number from root to the leaf as the representation to the tool in that leaf. The example in Appendix A may be assigned a number longer than five digits, such as 0 1 2 2 3 3 3.
Appendix C Constrained Beam Search Detail
@@ Appendix D Retry Mechanism
Appendix E Ablation

### v3

@@ Abstract
As large language models (LLMs) advance, their inability to autonomously execute tasks by directly interacting with external tools remains a critical limitation. Traditional methods rely on inputting tool descriptions as context, which is constrained by context length and requires separate, often inefficient, retrieval mechanisms. We introduce ToolGen, a paradigm shift that integrates tool knowledge directly into the LLM’s parameters by representing each tool as a unique token. This enables the LLM to generate tool calls and arguments as part of its next token prediction capabilities, seamlessly blending tool invocation with language generation. Our framework allows the LLM to access and utilize a vast amount of tools with no additional retrieval step, significantly enhancing both performance and scalability. Experimental results with over 47,000 tools show that ToolGen not only achieves superior results in both tool retrieval and autonomous task completion but also sets the stage for a new era of AI agents that can adapt to tools across diverse domains. By fundamentally transforming tool retrieval into a generative process, ToolGen paves the way for more versatile, efficient, and autonomous AI systems. ToolGen enables end-to-end tool learning and opens opportunities for integration with other advanced techniques such as chain-of-thought and reinforcement learning, thereby expanding the practical capabilities of LLMs [^1] Data and code are available at https://github.com/Reason-Wang/ToolGen.
@@ 1 Introduction
Large language models (LLMs) have demonstrated impressive capabilities as interactive systems, adept at processing external inputs, executing actions, and autonomously completing tasks (Gravitas, 2023; Qin et al., 2023; Yao et al., 2023; Shinn et al., 2023; Wu et al., 2024a; Liu et al., 2024; Wang et al., 2024b; c). Among the various methods enabling LLMs to interact with the world, tool calling via APIs has emerged as one of the most common and effective approaches. However, as the number of tools grows into the tens of thousands, existing methods for tool retrieval and execution struggle to scale efficiently.
@@ 2.1 Tool Retrieval
Recently, generative retrieval has emerged as a promising new paradigm, wherein models directly generate relevant document identifiers rather than relying on traditional retrieval mechanisms (Wang et al., 2022; Sun et al., 2023b; Kishore et al., 2023b; Mehta et al., 2023b; Chen et al., 2023c). Motivated by this, ToolGen represents each tool as a unique token, allowing tool retrieval and calling to be framed as a generation task. Beyond simplifying retrieval, this design integrates smoothly with other LLM and LLM-based agent features like chain-of-thought reasoning (Wei et al., 2023) and ReAct (Yao et al., 2023). By consolidating retrieval and task execution into a single LLM agent, it reduces latency and computational overhead, leading to more efficient and effective task completion.
@@ 2.2 LLM-Agents with Tool Calling
LLMs have shown strong potential in mastering tools for various tasks. However, most existing works focus on a limited set of actions (Chen et al., 2023a; Zeng et al., 2023; Yin et al., 2024; Wang et al., 2024a). For instance, Toolformer (Schick et al., 2023) fine-tunes GPT-J to handle just five tools, such as calculators. While effective for narrow tasks, this approach struggles in real-world scenarios with vast action spaces. ToolBench (Qin et al., 2023) expands the scope by introducing over 16,000 tools, highlighting the challenge of tool selection in complex environments.
To perform tool selection, current methods often use a retriever-generator pipeline, where relevant tools are retrieved and then utilized by the LLM (Patil et al., 2023; Qin et al., 2023). In addition, TPTU (Ruan et al., ) proposes a structured framework for LLM agents and evaluates their task planning and tool usage abilities. Furthermore, TPTU-v2 (Kong et al., ; Kong et al., 2024) builds an LLM Finetuner to enhance agent performance with curated datasets and a demo selector to select relevant demonstrations. They set a flexible and superior paradigm compared to traditional retrieval-based paradigm. However, pipelined approaches face two major issues: error propagation from the retrieval step and the inability of LLMs to fully understand and use tools via simple prompting.
@@
Our approach differs from ToolkenGPT in several ways. First, we focus on real-world tools that require flexible parameters for complex tasks (e.g., YouTube channel search), while ToolkenGPT is limited to simpler tools with fewer inputs (e.g., math functions with two numbers). Additionally, ToolkenGPT relies on few-shot prompting, whereas ToolGen incorporates tool knowledge directly into the LLM through full-parameter fine-tuning, enabling the model to retrieve and execute tasks autonomously. Finally, our experiments involve a much larger tool set—47,000 tools compared to ToolkenGPT’s 13–300. Detailed comparison and other related work can be found in Section A.
@@ 3.6 Inference
During inference, the LLM may generate action tokens outside the predefined tool token set. To prevent this, we designed a constrained beam search generation that restricts the output tokens to the tool token set. We applied this constrained beam search for both tool retrieval, where the model selects tools based on queries, and the end-to-end agent system, significantly reducing hallucination during the action generation step. A detailed analysis can be found in Section 5.4. The implementation details can be found in Appendix E.
@@ 4.1 Experimental Setup
Our experiments are based on ToolBench, a real-world tool benchmark containing more 16k tool collections, each containing several APIs, resulting in a total of 47k unique APIs. Each API is documented with a dictionary, containing the name, description, and parameters for calling the API. A real example is shown in Appendix C. We take each API as an action and map it to a token. Our retrieval and end-to-end agent-tuning data are converted from the original data in ToolBench. Details can be found in Appendix K. Although each tool may consist of multiple APIs, for simplicity, we refer to each API as a tool in this paper.
@@
  • Long-Context LLMs: We concatenate tools into a long prompt to gpt-4o, and prompt it to choose from the pool. Limit by context length, we cannot input all 47k tools, so we use 2k tools with ground truth tools included.
@@ 4.3 Indexing Method Comparison
Figure 3: The distribution of the number of subtokens per tool varies across different indexing methods.
The implementation details are described in Appendix D.
First, we conducted an analysis of the number of subtokens required to represent each tool for the different methods, as shown in Figure 3. Atomic indexing ensures each tool to be a single token, while numeric indexing encodes tools into N tokens for tools numbered in $(10^{N-1},10^{N}]$. In contrast, both semantic indexing and hierarchical indexing produce a variable number of subtokens, with semantic indexing having more outliers with significantly longer sequences. The figure highlights the superiority of atomic indexing, where each tool is represented by a single token, whereas other methods require multiple tokens. This efficiency allows ToolGen to reduce the number of generation tokens and inference time in both the retrieval and agent scenarios.
Next, we examined the effectiveness of different indexing methods. As shown in Table 2, semantic indexing demonstrates the best retrieval performance across various metrics and scenarios, while atomic indexing closely follows in many cases. We attribute this to the fact that semantic indexing aligns better with the pretraining data of LLMs. However, this advantage diminishes as the training data and type increase. For example, in Section 5.3, we show that atomic indexing achieves better end-to-end results. We also show that combining constrained beam search with semantic indexing will cause biased tool usage, which is detailed in Section E.2.
Taking all these factors into account, we choose atomic indexing for ToolGen tool virtualization.
@@ 4.4 Ablation
We perform an ablation study to assess the impact of different training stages of ToolGen, as shown in Table 3. The results indicate that retrieval training is the crucial factor for tool retrieval performance, as it directly aligns with the retrieval task where inputs are queries and outputs are tool tokens. Removing tool memorization leads to a minor performance drop although it plays a role in improving generalization, which we will discuss further in Appendix J. Similarly, constrained beam search, while not a major contributor to retrieval task, helps prevent hallucinations, making it useful for end-to-end agent tasks, see Section 5.4.
@@ 5.1 Experimental Setup
We make several modifications to the trajectory data from ToolBench to fit it into ToolGen framework. For example, as ToolGen does not require explicit selection of related tools as input, we remove this information in the system prompt. Further details are provided in Appendix K. Following this, we fine-tune the retrieval model using the reformatted data, resulting in an end-to-end ToolGen agent.
@@
GPT-3.5: We use gpt-3.5-turbo-0613 as one of our baselines. The implementation is the same as used in StableToolBench (Guo et al., 2024), where the tool calling capability of GPT-3.5 is used to form a tool agent. ToolLlama-2: Qin et al. (2023) introduced ToolLlama-2 by fine-tuning Llama-2 (Touvron et al., 2023) model on ToolBench data. ToolLlama-3: To ensure a fair comparison, we fine-tuned Llama-3, the same base model used in ToolGen, on the ToolBench dataset, creating the ToolLlama-3 baseline. In the rest of this paper, we refer to ToolLlama-3 as ToolLlama to distinguish it from ToolLlama-2.
@@
w/ Ground Truth Tools (G.T.) Following Qin et al. (2023), we define ground truth tools for a query as those selected by ChatGPT. For ToolLlama, we directly input the ground truth tools in the prompt, consistent with its training data format. For ToolGen, which is not trained on data with pre-selected tools, we add a prefix during the planning phase: I am using the following tools: [tool tokens], where [tool tokens] are virtual tokens corresponding to the ground-truth tools. w/ Retriever In the end-to-end experiments, we use a retrieval-based setting. For baselines, we use the tools retrieved by ToolRetriever as the relevant tools. In contrast, ToolGen generates tool tokens directly, so no retriever is used.
@@
Additionally, we introduce a retry mechanism for all models to prevent early termination, the details are introduced in Section G. Specifically, if a model generates a response containing give up or I’m sorry, we prompt the model to regenerate the response with a higher temperature.
@@ 5.2 Results
Table 4 presents the end-to-end evaluation performance of various models in two settings: using Ground Truth Tools (G.T.) and a Retriever (R.). In the G.T. setting, ToolGen achieves the best average SoPR score of 54.19, outperforming GPT-3.5 and ToolLlama, with SoWR also highest for ToolGen at 49.70. In the Retriever setting, ToolGen maintains its lead with an average SoPR of 53.28 and SoWR of 51.51. ToolLlama shows competitive performance, surpassing ToolGen on some individual instances. An ablation study of end-to-end ToolGen is provided in Appendix K.
@@ 6 Conclusions
Appendix A More Related Work
Previous work include Toolformer and ToolkenGPT, already employed vocabulary expansion for tool learning. The main difference between our work and the others is: previous studies primarily demonstrate that through SFT (in Toolformer) or adding new tool tokens with pre-computed embeddings (in ToolKenGPT), LLMs can learn to use a very small number of tools. However, in real-world tool-calling (agent) scenarios, previous methods require listing available tools in the prompt, which greatly limits their practical use. Examples can be seen in Figure 5.
Figure 5: Real examples from ToolkenGPT, Toolformer, and ToolGen (ours). Both ToolkenGPT and Toolformer describe tools available in the prompt, while ToolGen does not require tools been mentioned in its prompt.
Appendix B Tool Extension and Maintenance
In ToolGen and other generative retrieval systems, tools or documents are embedded into the model’s parameters. Therefore, how to add and maintenance new tools/documents become challenging. For ToolGen, it not only generates the proper tool, but also fetches the documentation for that tool. If there are minor changes that the tool usage scenarios keep the same (e.g. small parameter changes), it can still generate the tool and rely on the fetched documentation to do further tasks.
For vast changes that the usage scenarios are different or adding totally new tools, we admit that ToolGen is not able to utilize these tools. However, this inefficiency exists and is persistent for generative retrieval systems Sun et al. (2023a); Chen et al. (2023b); Mehta et al. (2023a). Current methods to adapt these changes include continual training and constrained optimization (Mehta et al., 2023a; Kishore et al., 2023a), which we believe could also be applied to ToolGen to alleviate the above challenges.
Despite that ToolGen is inefficient of adopting to new tools, its unified design lead to unique advantages such as easy integration with Chain-of-Thought (Wei et al., 2023), Reinforcement Learning with Human Feedback (Ouyang et al., 2022), and inference time scaling (Brown et al., 2024; Snell et al., 2024; Wu et al., ). We leave the problem of maintaining and adding tools to future work.
Appendix C Real Tool Example
@@
Appendix D Tool Virtualization Implementation
@@
indexing is the method we use in ToolGen. Compared to other methods, it takes a single token as a tool and does not hallucinate to nonexistent tools. We use <<tool name&&api name>> to combine the tool name and api name to form a single token. For example, for the example in Appendix C, the resulting token is <<Youtube Hub&&Get Video Details>>.
@@
indexing maps each tool to the name used in ToolBench, which is also a combination between tool name and API name. However, the name can be tokenized into multiple tokens so that the model can perceive its semantic meanings. For the example in Appendix C, the resulted mapping is get_video_details_for_youtube_hub.
@@
indexing maps each tool to a unique number. We first get a list of all tools, with a length about 47,000. For all tools, we use a five digit number separated by space to represent the tool. If the example in Appendix C is the 128th element in the list, we use 0 0 0 1 2 8 to represent the tool. Since Llama-3 tokenizer encodes each number separately, numeric indexing will lead to tool tokens with same number of sub-tokens.
@@
also maps each tool into a number. Different from Numeric indexing, we inject structure information into the tool representation by iterative clustering. During each iteration, we cluster tools into ten clusters, where each cluster is assigned a number from 0 to 9. For each cluster, we repeat this clustering process until there is only one tool in the cluster. These steps form a clustering tree. We take the number from root to the leaf as the representation to the tool in that leaf. The example in Appendix C may be assigned a number longer than five digits, such as 0 1 2 2 3 3 3.
Appendix E Constrained Beam Search
E.1 Implementation
@@
E.2 Bias Analysis
For semantic indexing, the prevalent constrained beam search introduces bias toward tools with more subtokens after tokenization. Traditionally, beam search retrieves the top-k decoding sequences at each step, and the sequence probability is computed by multiplying the probabilities of each token (given previous tokens) in the sequence and then averaging by the token count. Consider the following example with two tools:
  • ToolA: get_music_from_us $\rightarrow$ [get, music, from, usa]
  • ToolB: get_music_from_spain_black_singer... (a long-tail name) $\rightarrow$ [get, music, from, spain, black, singer, …]
After tokenization, the first three tokens are identical. For the fourth token, suppose usa has a probability of 0.7, and spain has a probability of 0.3. However, after spain, ToolB has a long tail of tokens with no alternatives, resulting in all subsequent tokens having a probability of 1.
How should the best tool for decoding be determined in this scenario? Using the traditional method, ToolB’s sequence probability increases as its unique number of tokens grows with each time step. This results in that tools with more unique tokens will have a higher probability to be retrieved.
As shown above, semantic tool name encoding tends to produce many long-name tools, and hence such bias becomes severe. However, there is no common solution to this type of bias. Note that this problem also exists in constrained natural language decoding. Still, because language candidates are typically very large, and constraints are usually associated with ban list of tokens or words, this issue is not usually considered.
Based on the above observations, we made the hallucination comparison in paper, with a setting of non constrained beam search (to avoid length bias) for other encoding methods. For atomic encoding, hallucination and bias do not exist even with constrained decoding (not beam search) because each tool is represented by a single token, ensuring unbiased and deterministic decoding.
Appendix F Integrate Instruction-Following Data
To ensure the model’s general ability is not lost after tool-specific task training, we incorporated general instruction-following data OpenHermes-2.5 into each training stage of ToolGen. [^2] https://huggingface.co/datasets/teknium/OpenHermes-2.5 We used a 50:50 ratio of our tool data and instruction-following data, resulting in a model called ToolGen-Instruct, indicating its capability to follow general instructions.
We evaluate the models general ability using 6 widely used LLM evaluation benchmarks, including ARC-easy, ARC-challenge, Commonsense QA, Hellaswag, Winograde, and GSM8K, all with 3-shot in-context examples. Results are shown in Table 6.
Table 6: 3-shot evaluation results across different NLP benchmark tasks, including ARC Challenge, ARC Easy, Commonsense QA, Hellaswag, Winograde, and GSM8K. The average performance (AVG.) is calculated as the mean across all tasks.
Method ARC-C ARC-E CSQA Hellaswag Winograde GSM8K AVG.
Llama-3 50.34 80.09 69.37 60.13 73.32 49.28 63.76
Llama-3-Inst 57.16 85.31 78.05 58.69 76.24 74.45 71.65
ToolGen 20.14 33.88 19.66 31.61 54.62 1.74 26.94
ToolGen-Inst 51.62 82.49 78.79 56.33 73.16 62.02 67.40
From the table, we can see that ToolGen’s general capability is limited. The original Llama 3 got an average score of 63.76, while ToolGen obtained almost random results with a score of 26.94, showing a significant gap. However, ToolGen-Instruct shows significant improvement, resulting in a 67.4 average score.
To check ToolGen-Instruct’s tool learning results, we conduct the end-to-end evaluation, see Table 7. Surprisingly, we find that its average performance improved by 3-5 percentage points compared to ToolGen. We did a manual comparison and found that ToolGen-Instruct performs better on final output summarization and provides more positive responses to the user after tool calling. From this observation, we conclude that general-purpose training can help ToolGen provide more helpful responses. We leave further exploration to future work.
Table 7: End-to-end agent evaluation of ToolGen and ToolGen-Instruct
Method SoPR SoWR
I1 Inst. I2 Inst. I3 Inst. AVG. I1 Inst. I2 Inst. I3 Inst. AVG.
ToolGen 56.13 52.20 47.54 53.28 50.92 62.26 34.42 51.54
ToolGen-Inst 63.09 55.03 54.10 58.84 51.53 60.38 50.81 54.24
Appendix G Retry Mechanism
Appendix H ToolGen for Different Sizes of LLMs
We also investigate how ToolGen’s performance changes as the sizes of base models change. Llama-3 series of models are not suitable as we can only use the 8B model and the 70B is too large for us. Therefore, we select Qwen2.5 (Team, 2024) with sizes of 1.5B, 3B, 7B, and 14B. As shown in Figure 8, models with larger sizes achieve better performance in tool retrieval and agent tasks. When the model size reaches 7B, the performance tends to plateau. And scaling it to 14B does not necessarily improve performance. While for generalization, larger models do not show better generalization capability.
Figure 8: Performance and generalization of ToolGen with different sizes of LLMs as the base model. For tool retrieval, the performance is calculated as the average score of I1, I2, and I3 domains. For end-to-end and generalization evaluation, it is based on the average score of unseen instructions and tools respectively.
Appendix I Ablation